OpenStudy (anonymous):

x: 0, 1, 2, 3, 4, 5, 6, 7, 8, 9. y: 5, 10, 20, 40, 80, 160, 320, 640, 1280, 2560. Create an equation to represent y: y = ___________

OpenStudy (bahrom7893):

x = 0 y = 5 Doesn't tell us much It might be y = x+5, but we can't be sure yet, so let's continue.

OpenStudy (bahrom7893):

x = 1 y = 10 Well now we can confirm that y is not x+5, since 1+5=6, not 10. So we scratch our original equation.

OpenStudy (anonymous):

that might help if it turns out to be a quadratic

OpenStudy (bahrom7893):

Thanks, I was getting stuck lol

OpenStudy (bahrom7893):

It looks like it's \[2^x*5\] so far. Let's check..

OpenStudy (bahrom7893):

\[2^0*5=1*5=5\]\[2^1*5=2*5=10\]

OpenStudy (bahrom7893):

So far so good.. Let's try the next one: x=2; y = 20: \[y=2^x*5=>2^2*5=4*5=20\]And the next one: x=3; y=40: \[y=2^x*5=>2^3*5=8*5=40\]

OpenStudy (bahrom7893):

\[2^4*5=16*5=80\]

OpenStudy (bahrom7893):

etc. Just make sure they all follow this pattern, but I think the equation is: \[y=2^x*5\]

OpenStudy (anonymous):

Thank you!

OpenStudy (bahrom7893):

You still have to make sure it works for all the other numbers.

OpenStudy (anonymous):

It does.

OpenStudy (bahrom7893):

Good, then that's the answer.